OpenStudy (anonymous):

PLEASE HELP!!!! using the following system of inequalities find the maximum value of f(x,y)=3x+8y x≥0 y≥0 3x+2y≤18 6x+7y≤42

Nnesha (nnesha):

you need to solve last two equations for y or just use desmos calculator shade the region

OpenStudy (anonymous):

y≤-3/2x+9 y≤-6/7x+6

OpenStudy (anonymous):

$$x\le 0\\y\le 0\\y\le -\frac32 x+9\\y\le -\frac67 x+6$$